OpenStudy (konradzuse):

integral of e^3x sec(e^3x) tan(e^3x)dx

OpenStudy (konradzuse):

u = sec(e^3x) du =e^3x sec(e^3x) tan(e^3x dx? \[\int\limits du\]?? = u + c = sec(e^3x) +c?

OpenStudy (turingtest):

du=3e^3x sec(e^3x) tan(e^3x) dx but other than that, yes

OpenStudy (konradzuse):

\[3 \int\limits du\]??

OpenStudy (konradzuse):

err actually...

OpenStudy (konradzuse):

\[1/3 \int\limits du\]

OpenStudy (turingtest):

yep

OpenStudy (konradzuse):

I guess I forgot the chain rule oopsies :P

OpenStudy (turingtest):

you did well to leave it at du most would have tried to integrate u

OpenStudy (konradzuse):

\[1/3 \sec(e^{3x}) +c\]

OpenStudy (turingtest):

correct :)

OpenStudy (konradzuse):

Yeah I was first thinking "Wait what about u?" Then I checked the front of the book to which it stated int du = u+c so I figured I'd go with that and make sure :D.

OpenStudy (konradzuse):

btw for the last question tan(arcsec(x/8)) I wrote out y = arcsec(x/8) then sec y = (8/x) then tan y = that big fraction? Is that correct how to write it out?

OpenStudy (konradzuse):

That's what the book says, so hopefully :P.

OpenStudy (turingtest):

sure, that's how I wrote it out with the picture next to it, and y labeled, it makes more senseI guess I would prefer to call the angle u instead of y, so I'd write\[\sec^{-1}u=\frac x8\implies u=\sec\frac x8\implies\tan u=\frac{\sqrt{x^2-8^2}}8\]